find the slope 1/2x+6y=12
Answer by HyperBrain(694)

If the line is written in the form y=mx+b, then m is the slope.
1/2x+6y=12
Multiply 2 to both sides
x+12y=24
Subtract x from both sides
12y=-x+24
Divide both sides by 12.
y=-x/12 + 2
Now, this is of the form y=mx + b. So the slope which is equal to m is -1/12